OpenStudy (anonymous):

if f(x)=x^2-2x+4, find f(-3x)

OpenStudy (hba):

Just plug in -3x in place of x :)

OpenStudy (hba):

@AutumnMaybury Start working :)

OpenStudy (anonymous):

answer is one?

OpenStudy (hba):

@AutumnMaybury What ?

OpenStudy (anonymous):

i got one for the answer

OpenStudy (hba):

Please show me your working @AutumnMaybury

OpenStudy (anonymous):

f(x)=-3^2-2(-3)+4

OpenStudy (hba):

Its not -3,It's -3x

OpenStudy (hba):

See the question you provided us

OpenStudy (anonymous):

but you said put -3 in for x..... *confused*

OpenStudy (hba):

I said put '-3x' in place of x

OpenStudy (anonymous):

ok sorry rewriting now

OpenStudy (hba):

Ok :)

OpenStudy (anonymous):

3x+4

OpenStudy (hba):

@AutumnMaybury I want you to show me your working please :)

OpenStudy (anonymous):



OpenStudy (hba):

Well lemme show you something \[\huge\ (-3x)^2=(-3)^2(x)^2=9x^2\]

OpenStudy (anonymous):

is that the answer?? please explain *confused* :(

OpenStudy (hba):

It would be x^2 not x

OpenStudy (hba):

I mean to say the first one my bad.

OpenStudy (hba):

f(x)=9x^2-6x+4

OpenStudy (anonymous):

but 9x-6x does equal 3x-4

OpenStudy (hba):

It would be 9x^2 not 9x.

OpenStudy (anonymous):

is the 9 the only one squareD? the 6x isnt?

OpenStudy (hba):

No.

OpenStudy (hba):

The 6x is not.

OpenStudy (hba):

I have showed you above in latex.

OpenStudy (anonymous):

ok thank you

OpenStudy (hba):